Assume Z is standard normal. Find the following values

(A) P(−0.43 ≤ Z ≤ 0.43)

(B) P(Z ≥ 1.54)
Answer #1

Solution:

A)

P(-0.43 < z < 0.43)

= P(z < -0.43) - P(z < 0.43)

= 0.6664 - 0.3336 Using standard normal table,  

Probability = 0.3328

B)

P(Z ≥ 1.54) = 1 - P(Z < 1.54 )

= 1 - 0.9352   Using standard normal table,  

Probability = 0.0618